OpenStudy (anonymous):

The reception for a radio station is given by the equation (x+1)^2/225 + (y+3)^2/144=1 A straight road passes through his area with an equation of 3x=y Where x and y are measured in miles in both questions. A car enters the the reception area at point a and exits at point b, as shown in the graph below, traveling at 30 mph for this entire distance. For how much time can the car receive this station.? Show all work

OpenStudy (anonymous):

the graph is backwards

OpenStudy (mathstudent55):

The line in the graph cannot have the equation you wrote, 3x = y.

OpenStudy (anonymous):

he told us to use substitution to fine the two (x,y) coordinates and then use the distance formula

OpenStudy (baru):

@mathstudent55 the camera image is mirrored, i guess the left side is the positive x axis

OpenStudy (anonymous):

yea sorry about that

OpenStudy (mathstudent55):

Your line in the graph is less than 45 up from the horizontal. That means the slope is between 0 and 1. Your equation 3x = y is the same as y = 3x. With a slope of 3, the line would be much more than 45 degrees up fro the horizontal. Perhaps you mean the line to be 3y = x.

OpenStudy (baru):

in any case, we do not need a figure

OpenStudy (baru):

the curve mentioned is an ellipse, not centred at the origin, so we can disregard the figure

OpenStudy (mathstudent55):

I understand the picture is a mirror image, but unless you have the correct equation for the line, you will not be answering the question based on the line in the figure.

OpenStudy (baru):

@nadiainderjit disregard the diagram completely, you can just use the equations

OpenStudy (mathstudent55):

@nadiainderjit Can you show a picture of the problem exactly as it was given to you?

OpenStudy (baru):

to find a and b solve the two equations, substiute for y or x using the line eq into the curve eq

OpenStudy (baru):

substitute y=3x in the eq \(\frac{(x+1)^2}{225}+.....=...\)

OpenStudy (mathstudent55):

Thanks. You wrote ... (y + 3)^2, but it's ... (y - 3)^2, so be careful when solving the problem. I see that the equation is really 3x = y as you wrote. Then do what @baru suggested. Substitute 3x in for y in the equation of the ellipse. You will have a quadratic equation in x. Solve for x. You will get two solutions for x. Then insert each value of x in the equation 3x = y to get the corresponding y-coordinates. Each set of x and y is a point. Those two points are points A and B. Then use the distance formula to find the distance between A and B. Then use the speed to find how much time will take to travel that distance.

OpenStudy (anonymous):

i know i had to do that it was actually doing the math that confused me @mathstudent55

OpenStudy (mathstudent55):

If you need help, just show what you have and where you are stuck, and I'll try to help.

OpenStudy (anonymous):

i got as far as \[\frac{ x ^{2}+2x ^{2}+1}{ 225 }+\frac{ 9x ^{2}-18x+9} {144}=1\]

OpenStudy (mathstudent55):

Good. Now multiply both sides by the LCD, 144 * 225.

OpenStudy (mathstudent55):

Or you can just use the LCD to add the fractions on the left.

OpenStudy (anonymous):

so multiply by 32400

OpenStudy (mathstudent55):

You can just do this to get the LCD: \(\dfrac{ (x ^{2}+2x ^{2}+1)144}{ 225 \times 144 }+\dfrac{ (9x ^{2}-18x+9)225} {144 \times 225}=1\)

OpenStudy (anonymous):

\[\frac{ 144x ^{2}+288x+ 144}{32400}+\frac{ 2025x ^{2}-4050x+2025 }{ 32400 }=1\]

OpenStudy (mathstudent55):

Good. Now add the fractions using the common denominator.

OpenStudy (mathstudent55):

Do you get this? \(\dfrac{2169x^2 -3762x + 2169}{32400} = 1\)

OpenStudy (anonymous):

yes. Now what. Do i multiply by 32400 on both sides to get \[2169x ^{2}-3762x+2169=32400\]

OpenStudy (anonymous):

then subtract to get \[2169x ^{2}-3762x-30231\]

OpenStudy (baru):

the numbers are big, so lets try to simply as far as possible. try dividing throughout by 9

OpenStudy (anonymous):

\[\left( 241x ^{2}-418x-33359 \right)\] @baru

OpenStudy (baru):

*3359 (u hav put an extra 3) do you know the quadratic formula?

OpenStudy (anonymous):

yea \[-b \pm \sqrt{ \frac{ b2-4ac }{ 2a }}\]

OpenStudy (mathstudent55):

So far you are correct. Now use the quadratic formula to find 2 solutions for x.

OpenStudy (mathstudent55):

\( 241x ^{2}-418x-3359 = 0\)

OpenStudy (baru):

*2a of the denominator is outside the square root, and it divides 'b' also

OpenStudy (mathstudent55):

\(\large x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (anonymous):

(4.7, ) (-2.96, )

OpenStudy (anonymous):

(4.7, 14.1 ) (-2.96, -8.88)

OpenStudy (anonymous):

using the distance formula would the final distance be 24.25 miles

OpenStudy (mathstudent55):

Your points are correct.

OpenStudy (mathstudent55):

Your distance is correct.

OpenStudy (mathstudent55):

Now just find the time using the given speed.

OpenStudy (anonymous):

how?

OpenStudy (baru):

speed=\( \frac{distance}{time}\)

OpenStudy (baru):

speed is given. you just found the distance. substitute these

OpenStudy (anonymous):

24.5/30= TIME

OpenStudy (baru):

yep

OpenStudy (anonymous):

so about 49 minutes

OpenStudy (mathstudent55):

Correct. Great job!